# Beyond Text-to-SQL: Autonomous Research-Driven Database Exploration with DAR

### v1

@@ Abstract
Ostap Vykhopen, Viktoria Skorik, Maxim Tereschenko  and Veronika Solopova Correspondence: ostap.vykhopen@mantisanalytics.com
@@ 1 Introduction
Modern enterprises accumulate data at unprecedented scales, yet much of it becomes “dark data”, poorly indexed, rarely queried, and eventually lost schembera2019dark; heidorn2008shedding. Industry estimates suggest that over 90% of organizational data remains unused, representing a major missed opportunity george2023bringing. The bottleneck is no longer computation: modern warehouses answer complex queries in seconds, but require humans to notice opportunities, prepare data, and manually craft queries kuchnik2022plumber. As data collections grow, this effort scales poorly, leaving patterns undiscovered and cross-table relationships unexplored.
What enterprises need is autonomous, proactive exploration rather than purely reactive querying. AI-powered data exploration promises a shift from responding to isolated questions to running continuous research cycles that identify promising directions and follow-up analyses zheng2025automation.
@@ 2.1 Multi-Agent Systems and LLM Agents
Recent advances in LLM-based agents have demonstrated capability to perceive environments, make autonomous decisions, and orchestrate tools to accomplish complex tasks zhou2025autonomous. Multi-agent systems enhance the abilities of single agents by coordinating and collaborating, allowing agents to communicate and share problem-solving tasks across specialized roles.
SchemaAgent pioneered multi-agent frameworks for database schema generation, demonstrating how specialized agents can collaborate on database-related tasks through role assignment and error detection mechanisms wang2025text2schema. However, this system focuses on schema design rather than autonomous data exploration.
Beyond schema-focused agent frameworks, prior work on multi-agent decision-making typically assumes explicit task definitions and engineered reward structures. However, traditional multi-agent reinforcement learning methods need a lot of training and specific reward frameworks zhang2019multiagent, while autonomous exploration demands systems that proactively investigate without explicit reward signals.
@@ 2.2 Database Querying and Understanding
Text-to-SQL systems translate natural language queries into SQL commands, with modern LLM-based approaches demonstrating strong performance even on complex structures through retrieval-augmented generation, database metadata, and self-correction loops deng2022texttosql. However, text-to-SQL systems remain fundamentally reactive. Designed to answer explicit user questions rather than proactively exploring databases to uncover patterns, relationships, or insights. They excel at query generation but cannot formulate their own research questions or decide what to investigate.
Beyond query generation, current database agents can list tables, retrieve schemas, and execute queries hong2024nextgen; mohammadjafari2025natural. Yet their focus remains on structural understanding: examining metadata, tables, and columns. They rarely explore data content, quality characteristics, or statistical patterns across tables. Autonomous exploration requires investigating what the data contains, not just how it’s organized.
Recent advances in cloud platforms like BigQuery have introduced native generative AI capabilities support functions like AI.GENERATE, AI.GENERATE_BOOL, and AI.GENERATE_TABLE for structured data extraction, along with row-wise LLM inference for SQL-driven data manipulation google2025bigqueryai. BigQuery’s evolution toward an autonomous data-to-AI platform includes automated metadata generation, Gemini-assisted data preparation, and multimodal ObjectRef data types google2025bigqueryautonomous.
@@ 2.3 Knowledge Graphs and Data Exploration
GraphRAG methods show how LLM-generated knowledge graphs enhance retrieval-augmented generation through structured relationships and multi-hop reasoning hu2024grag. Enterprise knowledge graphs built through LLM-based entity extraction support advanced querying across heterogeneous sources kumar2025llmkg.
@@ 2.4 Research Gap
Existing text-to-SQL systems remain reactive to user prompts deng2022texttosql; hong2024nextgen, schema discovery focuses on structure alone marotta2018schema, knowledge graphs are typically decoupled from databases Mohamed2025KnowledgeGraphs; Tian2022KGReasoning, and database agents largely wait for human instructions Wang2024LLMAgentsSurvey. To the best of our knowledge, no prior system autonomously formulates research questions, explores data content, and generates insights without human query initiation.
@@ 4 Experimental setup
This solution is implemented using Google ADK, an open-source framework for agentic programming. LLM agents are powered by Gemini gemini2024family; gemini2024onepointfive. The code is available on Github [^1] https://github.com/MantisAnalytics/DAR.

### v2

@@ Abstract
Ostap Vykhopen, Viktoria Skorik, Maksym Tereshchenko  and Veronika Solopova Correspondence: ostap.vykhopen@mantisanalytics.com
@@ 1 Introduction
Modern enterprises accumulate data at unprecedented scales, yet much of it becomes “dark data”, poorly indexed, rarely queried, and eventually lost Schembera and Durán (2019); Heidorn (2008). Industry estimates suggest that over 90% of organizational data remains unused, representing a major missed opportunity George et al. (2023). The bottleneck is no longer computation: modern warehouses answer complex queries in seconds, but require humans to notice opportunities, prepare data, and manually craft queries Kuchnik et al. (2022). As data collections grow, this effort scales poorly, leaving patterns undiscovered and cross-table relationships unexplored.
What enterprises need is autonomous, proactive exploration rather than purely reactive querying. AI-powered data exploration promises a shift from responding to isolated questions to running continuous research cycles that identify promising directions and follow-up analyses Zheng et al. (2025).
@@ 2.1 Multi-Agent Systems and LLM Agents
Recent advances in LLM-based agents have demonstrated capability to perceive environments, make autonomous decisions, and orchestrate tools to accomplish complex tasks Zhou et al. (2025). Multi-agent systems enhance the abilities of single agents by coordinating and collaborating, allowing agents to communicate and share problem-solving tasks across specialized roles.
SchemaAgent pioneered multi-agent frameworks for database schema generation, demonstrating how specialized agents can collaborate on database-related tasks through role assignment and error detection mechanisms Wang et al. (2025). However, this system focuses on schema design rather than autonomous data exploration.
Beyond schema-focused agent frameworks, prior work on multi-agent decision-making typically assumes explicit task definitions and engineered reward structures. However, traditional multi-agent reinforcement learning methods need a lot of training and specific reward frameworks Zhang et al. (2019), while autonomous exploration demands systems that proactively investigate without explicit reward signals.
@@ 2.2 Database Querying and Understanding
Text-to-SQL systems translate natural language queries into SQL commands, with modern LLM-based approaches demonstrating strong performance even on complex structures through retrieval-augmented generation, database metadata, and self-correction loops Deng et al. (2022). However, text-to-SQL systems remain fundamentally reactive. Designed to answer explicit user questions rather than proactively exploring databases to uncover patterns, relationships, or insights. They excel at query generation but cannot formulate their own research questions or decide what to investigate.
Beyond query generation, current database agents can list tables, retrieve schemas, and execute queries Hong et al. (2024); Mohammadjafari et al. (2025). Yet their focus remains on structural understanding: examining metadata, tables, and columns. They rarely explore data content, quality characteristics, or statistical patterns across tables. Autonomous exploration requires investigating what the data contains, not just how it’s organized.
Recent advances in cloud platforms like BigQuery have introduced native generative AI capabilities support functions like AI.GENERATE, AI.GENERATE_BOOL, and AI.GENERATE_TABLE for structured data extraction, along with row-wise LLM inference for SQL-driven data manipulation Google Cloud (2025a). BigQuery’s evolution toward an autonomous data-to-AI platform includes automated metadata generation, Gemini-assisted data preparation, and multimodal ObjectRef data types Google Cloud (2025b).
@@ 2.3 Knowledge Graphs and Data Exploration
GraphRAG methods show how LLM-generated knowledge graphs enhance retrieval-augmented generation through structured relationships and multi-hop reasoning Hu et al. (2024). Enterprise knowledge graphs built through LLM-based entity extraction support advanced querying across heterogeneous sources Kumar et al. (2025).
@@ 2.4 Research Gap
Existing text-to-SQL systems remain reactive to user prompts Deng et al. (2022); Hong et al. (2024), schema discovery focuses on structure alone Marotta (2018), knowledge graphs are typically decoupled from databases Mohamed et al. (2025); Tian et al. (2022), and database agents largely wait for human instructions Wang et al. (2024). To the best of our knowledge, no prior system autonomously formulates research questions, explores data content, and generates insights without human query initiation.
@@ 4 Experimental setup
This solution is implemented using Google ADK, an open-source framework for agentic programming. LLM agents are powered by Gemini Gemini Team (2024b, a). The code is available on Github [^1] https://github.com/MantisAnalytics/DAR.